hoe many line segments are determined by the n vertices of an n-gon? of these, how many are diagnols?
Answer by stanbon(75887)

hoe many line segments are determined by the n vertices of an n-gon? of these, how many are diagnols?
---
Each pair of vertices determines a line.
# of line segments determined = nC2 = (n(n-1))/2
-------------------------------------
# of sides + number of diagonals = (n(n-1))/2
---
number of diagonals = (n(n-1))/2 - n
=====================================
Cheers,
Stan H.